ClusterAlg.gif (577 bytes) The ATLAS Level-1 Calorimeter Trigger

[Home] [Architecture] [Meetings] [Participants] [Publications] [Software] [Search]

ATLAS-UK Level-1 Calorimeter Trigger Meeting

Thursday 18 June 2002 at RAL

Present: Bruce Barnett, Ian Brawn, Adam Davis, James Edwards, Eric Eisenhandler (chair), John Garvey, Norman Gee, Stephen Hillier, Murrough Landon, Gilles Mahout, David Mills, Ed Moyse, Viraj Perera, Richard Staley, Alan Watson


Click this side                               Click this side
for summaries                                for slides (pdf)
Hardware status
Cluster Processor Module and test cards.............Richard
Cluster Processor Module and CP chip tests...........Gilles
Common Merger Module....................................Ian
CP/JEP ROD prototype..................................Bruce
DSS and ROD firmware updates..........................James
TCM, VME Mount Module, GIO cards and CAN...............Adam
Stockholm, Mainz and Heidelberg status.................Tony
Schedule and milestone schedule........................Eric
Software status
Software status....................................Murrough
Calibration and calorimeter discussions at CERN......Norman
Simulation and physics
Trigger simulator and integration status.................Ed
Secondary RoIs.........................................Alan
Input signals
TileCal cables and input connection document.......Murrough
TileCal receiver specification.........................Eric
Coming meetings
ATLAS week, Stockholm meeting, T/DAQ workshop..........Eric
Amsterdam and LECC papers.......................John/Gilles

Any other business
Dates of next UK meetings

Hardware status

Cluster Processor Module and test cards - Richard (slides)

CPM: Module clock distribution is working. VME interface is functioning, with access to ID , Revision No, Control and Status registers. FLASH memory controllers are giving reliable R/W and Erase operations, and have been used to sucessfully configure all Serialiser FPGAs. Serialiser control registers are now accessable from VME.

The CAN microcontroller is now running cool; overheatingwas caused by an unsoldered pin giving intermittent connection with crystal causing chip to run in a strange mode. This fault mode also caused the microcontroller to corrupt the module's internal Geographical Address lines.

The CPM's FPGA configuration circuit is not ideal for development circuits or dealing with fault conditions. All the FPGA configuration 'DONE' signals have been 'wire-anded' as suggested in various Xilinx application notes. However, should one or more FPGA not be configured, then this line will stay inactive even though all the other FPGAs have released the line. The configuration process will stall until this line goes active. However , the default configuration behaviour can be modified and the FPGAs made to 'wake-up' before releasing DONE. The next revision of CPM will condition the individual DONE signals to isolate any unconfigured part. The present FPGA Configuration Controller sends a fixed number of bytes to the FPGAs, not relying on the DONE signal.

There is still no solution to the mechanical strengthening of the module.

The next task is to use the Serialiser to play back 160 Mbit/s data. This requires the TTCrx controller to be working, which sources the 'EnPlayback' signal. We will attempt to configure all remaining FPGAs, and then test the TTCrx interface, for which we need a TTCdec card.

Probingthe module inside a crate is very difficult, and one solution is to build a 'pizza-box' style test crate, which horizontally holds the module under test, VME-VME link, plus a VME-- backplane with a suitable power supply. There was some discussion on how one could have better access with two adjacent modules under test, and the feasibility of building extender boards.

Summary. Making steady progress.

Cluster Processor Module and CP chip tests - Gilles (slides)

The decision to download only one FPGA serialiser instead of all 20, has caused some problems due to the interconnection of the DONE signal between all of them. The FPGA toggles from its configuration state to its operational state once the DONE signal of all other FPGAs has also been pulled up. This problem has been by-passed by generating dummy Serialiser firmware with the option to not ask for the DONE signal to go high. Once the dummy firmware has been proven to work correctly in the one FPGA under test, the 19 others have been downloaded and then the serialiser file itself has been loaded.

First VME access showed that registers were readable and writable. The firmware ID seems to be set to 0, as was seen during local simulation test. The dual-port RAM inside the serialiser is not accessible yet, the En-PlayBack signal coming from the TTCrx controller is not generated correctly. The controller has to be modified.

In order to download FPGA binary files, a new HDMC part has been written. This part, for the CPM only, reads a bit file, swaps each byte, and loads them into a flash memory via a FIFO. On VME request, the file is then passed to the FPGA. More parts have also been written for future tests: DatasourceTV, which reads an external file of test vectors to be download inside a RAM, and ModuleI2Cregister, reading and writing to TTCrx registers via the I2C interface implemented in the TTCrx controller.

The next step will be to download data in the RAM, download the CP chip with the ScanPath configuration, and check the validity of the transfered data.

Common Merger Module - Ian (slides)

Further problems with the FPGA DLLs have been encountered. One DLL was losing lock due to a dirty input clock signal, caused by beating between the 40 MHz system clock and a 25 MHz clock supplied to the I2C controller. Both of these clocks are currently provided by crystal oscillators. The problem has been solved temporarily by removing the 25 MHz oscillator.

Firmware has been produced with a playback functionality slightly different to that described in the CMM specification. Using this firmware, data has been loaded into the Crate-FPGA playback RAM, transmitted via the real-time data path to the System FPGA, and read from the System-FPGA playback RAM successfully. Firmware has also been produced for the FLASH-controller CPLD that will load the System FPGA with the contents of the FLASH RAM. This firmware has been simulated successfully but has yet to be tested in hardware.

In discussion, Norman pointed out the importance of using the TTC rather than a crystal as clock source.

CP/JEP ROD prototype - Bruce (slides)

Bruce presented an update on software and firmware status of items concerning the CP/JEP ROD. On the software front, code has been moved into repositories (CMT in the case of the moduleServices) to make public access possible. Integration work (towards a model comparable to that to be used in the slice tests) in conjunction with Murrough has incorporated the online database into the CP Rod test programme. A number of items here are pending, in particular the restructuring of test code into a "looper/kicker" pair of processes, and later incorporation of ROS.

Concerning firmware, much new DSS firmware is available (see James' presentation below) and will need to be tested in conjunction with the ROD.

Bruce outlined plans to set up a test-bench for use by others in pre-slice tests of the sixnew RODS which have been produced.

DSS and ROD firmware updates - James (slides)

James discussed the contents of Project Change Order 001 for the DSS, which can be downloaded from here. He briefly explained the various modifications and added features required for future ROD and GIO tests. Some of the changes are to provide memory wraparound that will allow capture of data on level-1 accepts.

In discussion, Norman pointed out that someone should update the module specifications to reflect these changes.

TCM, VME Mount Module, GIO cards and CAN - Adam (slides)

TCM - It has been noted that there is a problem with the TCM latching data from the backplane when accessing other modules. A new firmware version of the "VME decoder" and "data-latch-Visualisers" will be issued as soon as the problem has been solved.

VMM - The VME Mount Module has the problem of identifying what reset is performed when the front panel switch is pushed, and also the geographical address LEDs displaying the complement of the address. There may be a fix to both these problems, further development is required.

GIO - A mistake has been made when placing the components on the board and as a result, all of the CMC connectors are 180 degrees out. An adapter card is being manufactured and will be here by the end of this week or beginning of next. However, the card has been powered up and we can connect to the board via JTAG and see both PROMs and FPGA, we can also program the FPGA. The board has not been boundary scanned as yet but it will be when the adapter card comes.

CAN - Adam has managed to establish CAN connection between the CMM and TCM using the backplane of the processor crate, and he has further developed the code such that a network of three microcontrollers has been set up utilising the remote request facility of the CAN protocol, one master and two slaves.

Bruce suggested that future versions of the TCM might have CAN status lights for each module. Norman suggested that we try to capture all comments of this nature so that future modules incorporate all of the good suggestions.

Stockholm, Mainz and Heidelberg status - Tony (notes)

In Tony's absence there was no formal talk. Eric and Norman briefly described the last monthly hardware phone conference, but since then the notes have appeared and the reader is referred to them.


Schedule and milestone schedule - Eric

Eric pointed out that a lot has changed. Our timetable has slipped,and the LHC and the ATLAS completion dates have also slipped. Now that we have some idea of the new LHC schedule, we should re-think and update our own schedule and milestones. Tony has made a start on the UK part of this (see previous meetings), but we need to finish the job for the entire project and feed it into the ATLAS monitoring system. We have agreed with Nick that a new schedule and milestones would be discussed and approved at Stockholm.

Software status

Software status- Murrough (slides)

There has been software progress on several fronts. The HDMC CVS repository has now been moved to RAL. Gilles has been developing new HDMC parts for the CPM. The module services packages have now been released as CMT packages, and development of the CMM module services has begun. Integration of the module services with the database has started, though there are still some outstanding problems. The database classes have been extended to cover all the UK modules.

Some progress has been reported from Stockholm on the JEM simulation, and work on moving the simulation into CMT has started. However more work is required to integrate the simulation with the database and the run control.

The software group should try to make a more detailed timetable, and also do some software training.

Calibration and calorimeter discussions at CERN - Norman (slides)

Norman reported on discussions between himself, Murrough and Thomas, and calorimeter calibration contacts Rupert Leitner and Bob Stanneck (TileCal), and Pascal Perrodo and Isabelle Wingerter (LAr) during the recent DIG training week.

The TileCal has a caesium source, a laser pulsing system, and a charge injection system. We would probably use the charge injection, which can generate 30–50k events from 1–800 GeV per calorimeter cell in two minutes. The TileCal group have not so far developed a run control strategy for this sytem, and are very flexible about working together in a combined run style.

The LAr has a charge injection system which runs at 10 kHz covering a range of amplitudes, gains, patterns, etc. The system is controlled by a dedicated hardware calibration board which uses a Busy handshake with the LAr RODs. There is no use of the online run control system planned to handle the runnung, and more work is needed to understand how the trigger can interact with and use this system.

It was agreed to prepare a (3-way) joint document to collect and define the different procedures. Norman prepared a skeleton which Thomas will populate for the Stockholm meeting, in part using information from Murrough's earlier draft calibration note. There will be an opportunity for everyone interested to contribute to the ideas.

Norman also mentioned new ideas for a possible combined run with detectors, TDAQ and trigger in 2004. His proposal to ask TDSG to discuss this was endorsed. In discussion, it was agreed that the ambition level of such a test has to be kept in line with what can be achieved, that many goals do not actually require beam, and that many cross-interface tests can be factorised out and tested beforehand.

Simulation and physics

Trigger simulator and integration status - Ed (slides)

Ed summarised the current status of TrigT1Calo simulation. The ROD is almost complete with only S-links left to implement. A new RoIDecoder class has been introduced which returns coordinates from RoI words. The energy trigger is started and work has begun integrating TrigT1Calo with TrigT1CTP.

The simulation is needed for Data Challenge 1, and so needs urgently to be validated.

Secondary RoIs - Alan (slides)

The distinction between "primary" and "secondary" RoIs was reviewed. In a typical trigger menu, some thresholds may be used only for creating secondary RoIs, i.e. will not contribute to the level-1 decision. Nevertheless, the multiplicities of objects passing such thresholds will be reported to the CTP (which will not use the information). This is not a problem provided we have sufficient thresholds, and this usage was included in estimates of the number of thresholds we would need. Should a future trigger menu demand more thresholds than we anticipated, it should be possible (with firmware and software changes) to provide non-trigger RoIs which do not use up part of the "hit multiplicity" traffic to the CTP. This option will be kept in reserve unless needed.

Input signals

TileCal cables and input connection document - Murrough (slides)

Cables - A meeting with TileCal people was recently held at CERN. They promised to provide more detailed information about signal cables and pinouts. Due to space problems, it is now suggested to combine the calorimeter and muon trigger signals on a single 16-pair cable, at least from the extended barrel, which would be split at a patch panel in USA15. This solution has the advantage that we could use the same type of cable as the LAr, and so is very attractive. It would also make the choice of whether we need a patch panel before the receivers, since in this option we would.

The TileCal group suggested that we bring some of our electronics to their 2003 test beam. This would be difficult for us in view of the slice test.

Cabling document - Some feedback and corrections on our spreadhseets have been received from Bill Cleland. These have yet to be incorporated. The text part of the document is not yet written.

TileCal receiver specification - Eric (slides)

We have been informed by Bill Cleland that the LAr receivers will be made at Pittsburgh; this has now been approved by the US DoE. The only sensible option for the TileCal receivers is to build them at Pittsburgh too, and to make them as similar to the LAr ones as possible. They now have people who are willing to do this, but to get approval (even though we pay) they need a signal-handling specification to be written. We are supposed to do that. In the absence of analogue electronics expertise, Eric has volunteered to try to do this, consulting the Rio group where necessary.

Coming meetings

ATLAS week, Stockholm meeting, T/DAQ workshop - Eric

Only two people from our trigger project will be attending the ATLAS overview week in Clermont-Ferrand. Sam has agreed to give a talk on our status, and has posted it on the web for comments.

Preparations for the Stockholm meeting on 4–6 July are going well. One big decision is to agree on a new milestone schedule (see above).

The TDAQ workshop is the week after Stockholm. Norman will be going. Although this will concentrate on the DAQ/HLT TDR preparations, there are always useful things as well as the opportunity to talk to non-TDAQ people at CERN.

Amsterdam and LECC papers - John/Gilles

John has prepared a draft paper for the Amsterdam conference, and will distribute it for comments. The talk will have to be combined with calorimeter papers by Arizona (FCAL) and the LAr. Our preferred choice of speaker is Pascal Perrodo, who has submitted the LAr paper.

Any other Business

There was no other business.

Next UK meeting

The next UK trigger meeting was tentatively set for Wednesday, 24 July at RAL.

Eric Eisenhandler, 26 June 2002

This page was last updated on 16 December, 2013
Comments and requests for clarification may be sent to